OpenStudy (loser66):

Someone explains me, please \(f(x) =x^{p-1}+x^{p-2}+.....+x^2+x+1\) \(f(x+1) = \sum_{s=0}^{p-1}\left(\begin{matrix}p\\s+1\end{matrix}\right) x^s\) How to get f(x+1) from f(x) as above?

OpenStudy (loser66):

@tkhunny

OpenStudy (tkhunny):

1 1 1 1 2 1 1 3 3 1 1 4 6 4 1 1 5 10 10 5 1 1 6 15 20 15 6 1 p = 2 -- f(x+1) = x+2 = [(x+1)^2 - 1]/x p = 3 -- f(x+1) = x^2 + 3x + 3 = [(x+1)^3 - 1]/x p = 4 -- f(x+1) = x^3 + 4x^2 + 6x + 4 = [(x+1)^4 - 1]/x p = 5 -- f(x+1) = x^4 + 5x^3 + 10x^2 + 10x + 5 = [(x+1)^5 - 1]/x Starting to seem like you can just collect the like terms. It might seem tedious, at first.

OpenStudy (loser66):

I know pascal, and I know how to expand the sum. But I don't see the link between f(x) and f(x+1) For example: if p =3, then f(x) = x^2+x+1 f(x+1) =(x+1)^2 +(x+1)+1= x^2+3x+3

OpenStudy (zarkon):

\[f(x)=\sum_{k=0}^{p-1}x^k\] \(\displaystyle f(x+1)=\sum_{k=0}^{p-1}(x+1)^k=\sum_{k=0}^{p-1}\sum_{s=0}^{k}{k\choose s}x^k\) [binomial expansion of \((x+1)^k\)] \[0\le s\le k\le p-1\] \[=\sum_{s=0}^{p-1}\sum_{k=s}^{p-1}{k\choose s}x^s\] \[=\sum_{s=0}^{p-1}x^s\sum_{k=s}^{p-1}{k\choose s}\] by classifying the s+1 subsets of {1,2,3,...,p} according to the last element s+1 we have... \[\sum_{k=s}^{p-1}{k\choose s}={p\choose s+1}\] and thus \[=\sum_{s=0}^{p-1}x^s\sum_{k=s}^{p-1}{k\choose s}=\sum_{s=0}^{p-1}x^s{p\choose s+1}=\sum_{s=0}^{p-1}{p\choose s+1}x^s\]

OpenStudy (loser66):

Thank you.

OpenStudy (zarkon):

Just noticed...the second line has a typo...should be \[\Large f(x+1)=\sum_{k=0}^{p-1}(x+1)^k=\sum_{k=0}^{p-1}\sum_{s=0}^{k}{k\choose s}x^{\color{red}{s}}\]

OpenStudy (tkhunny):

Rewriting like that is the only step I was seeing. Glad you knew where to go from there.